OpenStudy (anonymous):

Algebra 2 help please :)

OpenStudy (jessiegonzales):

Question?

OpenStudy (anonymous):

Sure

OpenStudy (imstuck):

What do you need for Algebra 2?

OpenStudy (anonymous):

What is the slope intercept form equation of the line that passes through (2,4) and (4,10) A. y=-3x-2 B. y=-3x+2 C. y=3x+2 D. y=3x-2

OpenStudy (imstuck):

First find the slope using the points they gave you. Do you know how to do that?

OpenStudy (anonymous):

y=mx+b? right

OpenStudy (imstuck):

No, that's the slope-intercept form of an equation.

OpenStudy (imstuck):

Slope formula is this:

OpenStudy (imstuck):

\[m=\frac{ y _{2} -y _{1}}{ x _{2}-x _{1} }\]

OpenStudy (anonymous):

no no the slope would be m=3? right?

OpenStudy (imstuck):

Using your points fill the formula in like this:

OpenStudy (anonymous):

sorry I got confused

OpenStudy (imstuck):

NP : )

OpenStudy (imstuck):

\[m=\frac{ 10-4 }{ 4-2 }=\frac{ 6 }{ 2 }=3\]

OpenStudy (anonymous):

So the slope would be m=3

OpenStudy (imstuck):

You were right...the slope is 3. Now use one of your points to write the equation. Use the point-slope formula. Do you know it?

OpenStudy (anonymous):

y=y1=m(x-x1) ?

OpenStudy (anonymous):

y-y1=m(x-x1)**

OpenStudy (imstuck):

Yes that's it. Now pick a point and fill in your y1 and x1, using the slope of 3 that you found. Can you do that?

OpenStudy (anonymous):

y-4=3(x-2) ??? @IMStuck

OpenStudy (imstuck):

So sorry, was doing calculus and acceleration!

OpenStudy (anonymous):

its okay! :)

OpenStudy (imstuck):

Yes that is the right equation using the point (2,4) and the slope of 3. Good!

OpenStudy (imstuck):

Now you just have to rewrite it using the slope-intercept form of the equation. Solve it for y. Can you do that?

OpenStudy (anonymous):

y=3x+2 ?

OpenStudy (imstuck):

Close...it's y = 3x-2. When you distribute the 3 into the parenthesis, you get 3x-6, so when you add the 4, you have a -2, not a +2.

OpenStudy (imstuck):

Do you see your error?

OpenStudy (anonymous):

Yes I see where I went wrong. I didnt put the negative on the 2

OpenStudy (anonymous):

Could you help me with a few more sorry if im bothering you

OpenStudy (anonymous):

Hey can help

OpenStudy (anonymous):

I mean can try

OpenStudy (anonymous):

Thank you :)

OpenStudy (anonymous):

Ok

OpenStudy (anonymous):

No problem

OpenStudy (anonymous):

Warning I am NOT GOOD AT MATH

OpenStudy (anonymous):

given the function f(x)=6(x+2)-3 solve for the inverse function when x=21 A.-45 B.0 C.2 D.63

OpenStudy (anonymous):

Lol its okay. Im not either

OpenStudy (anonymous):

what do not know just a sec ill look it up on a calculator :)

OpenStudy (anonymous):

Is that okay?

OpenStudy (anonymous):

Yes thats fine :)

OpenStudy (anonymous):

Here ill give you the links of the ones I'm going to use. http://www.wolframalpha.com/ and https://www.mathway.com/

OpenStudy (anonymous):

Thank you:)

OpenStudy (anonymous):

Your welcome click on this-> http://www.wolframalpha.com/input/?i=f%28x%29%3D6%28x%2B2%29-3

OpenStudy (anonymous):

And see if this is your answer

OpenStudy (anonymous):

No i didnt see the answer on there but its okay

OpenStudy (anonymous):

This should be the answer okay? (click on link.) https://www.mathway.com/ and type in the problem here is what I see f-1(x)=1/6x- 3/2 this sign / like with the 1/6 and the 3/2 these are fractions but type it in to make sure okay? :)